If the complement of an angle is 5 degrees less than one-sixth of its supplement, find the measure of the angle.
Answer by stanbon(48558)

If the complement of an angle is 5 degrees less than one-sixth of its supplement, find the measure of the angle.
------------------------------------------
Angle: x
Complement: 90-x
Supplement:180-x
-------------------------
Equation:
90-x = (1/6)(180-x)-5
---
90-x = 30 - (x/6) -5
65 = (5/6)x
x = 78
=====================
Cheers,
Stan H.